The length of a rectangle exceeds twice its width by 4 cm. The perimeter of the rectangle is 68 cm. Find the length and the width. (Hint: P=2l+2w)
Can you please help me with this question?

Answer by stanbon(75887)

The length of a rectangle exceeds twice its width by 4 cm. The perimeter of the rectangle is 68 cm. Find the length and the width. (Hint: P=2l+2w)
---------------------------
Perimeter = 2(L + W)
68 = 2((2W+4)+W)
34 = 3W + 4
30 = 3W
W = 10 cm (width)
L = 2W+4 = 2*10+4 = 24 cm (length)
=====================================
Cheers,
Stan H.